OpenStudy (anonymous):

Minimize F=x^2+y^2 with x+4y=68

OpenStudy (anonymous):

x+4y=68 is a constraint?

OpenStudy (anonymous):

Correct, yes it is

OpenStudy (anonymous):

G= Constraint equation \[\nabla F=\lambda \nabla G\]

OpenStudy (anonymous):

\[\nabla F={<2x,2y>}\] \[\nabla G={<1,4>}\]

OpenStudy (anonymous):

\[2x=\lambda*1\] \[2y=\lambda*4\] x+4y=68

OpenStudy (anonymous):

what does lambda mean?

OpenStudy (anonymous):

what does \[\lambda\] mean

OpenStudy (anonymous):

Just a constant, you can use C instead of \[\lambda\]

OpenStudy (anonymous):

Dude you are awesome, I appreciate it so much

OpenStudy (anonymous):

From top two equation , we can see that y=4x x+4y=68 x+16x=68 17x=68 x=4

OpenStudy (anonymous):

plug that back in to find y

OpenStudy (anonymous):

plug x=4 back into which equation?

OpenStudy (anonymous):

x+4y=68 4y=64 y=16 @ btw , what Calculus are you taking?

OpenStudy (anonymous):

Calc 2.... been absent so im playing catch up